if 120 of the 200 students are girls ,then what percent of the students are girls?
Found 2 solutions by mananth, richwmiller:
Answer by mananth(16946)

out of 200, 120 are girls
out of 100 how many will be girls
100*120 / 200
=60% are girls

Answer by richwmiller(17219)

120/200=60/100=60% girls
80/200=40/100=40% not girls